Cheryl is two more than two times as old as Henry. In 5 years the sum of their ages will be 78. How old is Cheryl now?
22
48
46
23

Answer by Cromlix(4381)

Hi there,
'
Cheryl is two more than two times as old as Henry
C = 2 + 2H.........(1)
In 5 years the sum of their ages will be 78.
C + 5 + H + 5 = 78........(20
Replace 'C' in Equation (20 with (2 + 2H)
from Equation (1)
2 + 2H + 5 + H + 5 = 78
Collect like terms:
H + 2H = 78 - 12
3H = 66
H = 22
Henry is currently 22 years old
Cheryl is currently 46 years old
Hope this helps :-)